# Deformations of pseudocharacters and Mazur’s finiteness condition

### v1

@@ 7. Completions at maximal ideals and deformation problems
Let $\kappa$ be a local field, which is an $\mathcal{O}$-algebra. Let $\mathcal{O}_{\kappa}$ be its ring of integers and let $k^{\prime}$ be the residue field. We always equip $\kappa$ with the topology induced by the valuation. Let $\mathcal{O}^{\prime}=\mathcal{O}\otimes_{W(k)}W(k^{\prime})$, where $W(k)$ is the Witt ring of $k$.
In this section we will study deformation theory of $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(\kappa)$, where $\Gamma$ is an arbitrary profinite group and $G$ is a generalised reductive group scheme over $\mathcal{O}$. If $\mathrm{char}(\kappa)=0$ then we let $\Lambda=\kappa$ equipped with its natural topology and we let $\Lambda_{0}=\mathcal{O}_{\kappa}$. If $\mathrm{char}(\kappa)=p$ then we let $\Lambda=\mathcal{O}\otimes_{W(k)}C$, where $C$ is a Cohen ring of $\kappa$, which means that $C$ is a complete DVR with uniformiser $p$ and residue field $\kappa$. If we choose an isomorphism $k^{\prime}\llbracket t\rrbracket\cong\mathcal{O}_{\kappa}$ of local $k^{\prime}$-algebras then $C$ can be chosen to be the $p$-adic completion of $W(k^{\prime})\llbracket t\rrbracket[1/t]$. There is a natural topology on $\Lambda$, such that the quotient topology on $\kappa$ induced via $\Lambda\twoheadrightarrow\kappa$ coincides with the valuation topology on $\kappa$. The topology on $\Lambda$ is Hausdorff, see [3, Section 3.5] for more details. We let $\Lambda_{0}=\mathcal{O}^{\prime}\llbracket t\rrbracket$. In both cases we let $\mathfrak{A}_{\Lambda}$ be the category of local artinian $\Lambda$-algebras with residue field $\kappa$. Any $A\in\mathfrak{A}_{\Lambda}$ is a finitely generated $\Lambda$-module, and hence carries a unique Hausdorff topology, which makes $A$ into a topological $\Lambda$-module.
For $A\in\mathfrak{A}_{\Lambda}$ we let $\operatorname{Def}_{\Theta}(A)$ be the set of $\widetilde{\Theta}\in\mathrm{cPC}^{\Gamma}_{G}(A)$ such that $\widetilde{\Theta}\otimes_{A}\kappa=\Theta$. It follows [^2] If $\mathrm{char}(\kappa)=p$ then $\Lambda$ is taken to be equal to $\kappa$ in [14]. However, the same argument goes through in our more general setting. from [14, Theorem 5.4] that the functor $\operatorname{Def}_{\Theta}:\mathfrak{A}_{\Lambda}\rightarrow\operatorname{Set}$ is pro-represented by a Hausdorff topological local $\Lambda$-algebra $R^{\mathrm{ps}}_{\Theta}$, which is a projective limit of rings in $\mathfrak{A}_{\Lambda}$. We let $\widetilde{\Theta}^{u}\in\mathrm{cPC}^{\Gamma}_{G}(R^{\mathrm{ps}}_{\Theta})$ be the the universal deformation of $\Theta$.
@@
By the reconstruction theorem [14, Theorem 3.7] there exists a continuous representation $\rho:\Gamma\rightarrow G(\overline{\kappa})$, such that $\Theta_{\rho}=\overline{\Theta}\otimes_{\kappa}\overline{\kappa}$. If $\kappa^{\prime}$ is a finite extension of $\kappa$ contained in $\overline{\kappa}$ then $G(\kappa^{\prime})$ is a closed subgroup of $G(\overline{\kappa})$ and we let $\Gamma_{\kappa^{\prime}}$ be its preimage in $\Gamma$. Then $\Gamma_{\kappa^{\prime}}$ is a closed and hence profinite subgroup of $\Gamma$. It follows from [6, Lemma 4.3] that there exists a finite extension $\kappa^{\prime\prime}$ of $\kappa^{\prime}$ contained in $\overline{\kappa}$ and $g\in G^{0}(\kappa^{\prime\prime})$ such that $g\rho(\Gamma_{\kappa^{\prime}})g^{-1}\subset G(\mathcal{O}_{\kappa^{\prime\prime}})$. This implies that $\Theta|_{\Gamma_{\kappa^{\prime}}}$ takes values in $\mathcal{O}_{\kappa^{\prime\prime}}$. Since $\Gamma=\bigcup_{\kappa^{\prime}}\Gamma_{\kappa^{\prime}}$ we deduce that $\Theta$ takes values in $\mathcal{O}_{\overline{\kappa}}\cap\kappa=\mathcal{O}_{\kappa}$. ∎
@@
Let $B$ be the $\mathfrak{q}$-adic completion of $\Lambda\otimes_{\mathcal{O}^{\prime}}R^{\mathrm{ps}}_{\overline{\Theta}}$. Then $B$ is a complete local noetherian $\Lambda$-algebra with residue field $\kappa$ by [4, Lemma 3.3.5], [3, Lemma 3.36]. Thus $B/\mathfrak{m}_{B}^{n}\in\mathfrak{A}_{\Lambda}$ for all $n\geq 1$ and $\widetilde{\Theta}_{B}:=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Theta}}}B$ is a deformation of $\Theta$ to $B$.
The ring $B$ pro-represents a subfunctor $\operatorname{Def}^{\prime}_{\Theta}$ of $\operatorname{Def}_{\Theta}$, such that $\widetilde{\Theta}\in\operatorname{Def}_{\Theta}(A)$ lies in $\operatorname{Def}^{\prime}_{\Theta}(A)$ if and only if $\widetilde{\Theta}$ takes values in a $\Lambda_{0}$-subalgebra $A_{0}$ of $A$, which is finite as a $\Lambda_{0}$-module. Indeed, in this case $\widetilde{\Theta}$ is a deformation of $\overline{\Theta}$ to $A_{0}$, and since $A_{0}$ is a limit of rings in $\mathfrak{A}_{\mathcal{O}^{\prime}}$ we obtain a morphism of local $\mathcal{O}^{\prime}$-algebras $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A_{0}$, which induces a morphism $B\rightarrow A$, such that $\widetilde{\Theta}=\widetilde{\Theta}_{B}\otimes_{B}A$. Moreover, $\widetilde{\Theta}_{B}\otimes_{B}B/\mathfrak{m}_{B}^{n}$ takes values in the image of $\Lambda_{0}\otimes_{\mathcal{O}^{\prime}}R^{\mathrm{ps}}_{\overline{\Theta}}$ in $B/\mathfrak{m}_{B}^{n}$, which is a finite $\Lambda_{0}$-module.
Since $\Gamma$ is finitely generated as a topological group, we may choose a dense discrete finitely generated subgroup $\Sigma$ of $\Gamma$. Since $\Sigma$ is dense the restriction map induces an injection $\operatorname{Def}_{\Theta}(A)\subset\mathrm{PC}^{\Sigma}_{G}(A)$. Let $A_{\infty}$ be the $\Lambda_{0}$-subalgebra of $A$ equal to the preimage of $\mathcal{O}_{\kappa}$ under $A\twoheadrightarrow\kappa$. Since $\mathfrak{m}_{A}$ nilpotent and a finite $\Lambda$-module, $A_{\infty}$ is a union of $\Lambda_{0}$-subalgebras $A_{n}$, $n\geq 1$, such that each $A_{n}$ is a finite $\Lambda_{0}$-module and the image of $A_{n}\rightarrow\kappa$ is equal to $\mathcal{O}_{\kappa}$. 7.1 implies that the restriction to $\Sigma$ identifies $\operatorname{Def}_{\Theta}(A)$ with a subset of $\mathrm{PC}^{\Sigma}_{G}(A_{\infty})$. Since $\Sigma$ is finitely generated, $\mathrm{PC}^{\Sigma}_{G}$ is represented by a finite type $\mathcal{O}$-algebra by [14, Proposition 3.21]. Hence, $\mathrm{PC}^{\Sigma}_{G}(A_{\infty})$ is the union of $\mathrm{PC}^{\Sigma}_{G}(A_{n})$ for $n\geq 1$. This implies that $\operatorname{Def}^{\prime}_{\Theta}(A)=\operatorname{Def}_{\Theta}(A)$, which implies the assertion. ∎
As in the previous section let $\overline{\rho}:\Gamma\rightarrow G(\bar{k})$ be a continuous $G$-semisimple representation such that $\Theta_{\overline{\rho}}=\overline{\Theta}$. Let $K$ be the kernel of $\overline{\rho}$ and let $Q$ be the quotient of $\Gamma$ defined in (12).
@@
It follows from 7.2 that $\operatorname{Ker}(\Theta^{u})\subset\operatorname{Ker}(\widetilde{\Theta}^{u})$. Since $K/\operatorname{Ker}(\Theta^{u})$ is pro-$p$ by 5.5, $K/\operatorname{Ker}(\widetilde{\Theta}^{u})$ is also pro-$p$. Thus $\Gamma/\operatorname{Ker}(\widetilde{\Theta}^{u})$ is a quotient of $Q$, which implies the equality of deformation rings. ∎
@@ 8. The rigid analytic space of $G$-pseudocharacters
Let $\mathrm{Rig}_{L}$ be the category of rigid analytic spaces over $L$. If $Y\in\mathrm{Rig}_{L}$ then its ring of global sections $\mathcal{O}(Y)$ carries a natural Hausdorff topology. If $Y=\operatorname{Sp}(A)$ is affinoid then the topology on $A$ is the Banach space topology. Let $\tilde{X}_{G}:\mathrm{Rig}_{L}^{\mathrm{op}}\rightarrow\operatorname{Set}$ be the functor that associates to every rigid analytic space $Y\in\mathrm{Rig}_{L}$ the set of continuous $G$-pseudocharacters $\mathrm{cPC}^{\Gamma}_{G}(\mathcal{O}(Y))$. We show that if $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then $\tilde{X}_{G}$ is represented by a rigid analytic space.
@@
Let $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(A)$, where $A$ is an affinoid $L$-algebra. Then $\Gamma/\operatorname{Ker}(\Theta)$ is an extension of a pro-$p$ group by a finite group.
@@
Let $\kappa_{i}$ be the residue field of $\mathfrak{m}_{i}$. Since $A$ is an affinoid algebra, $\kappa_{i}$ is a finite extension of $\mathbb{Q}_{p}$. Let $\Theta_{i}=\Theta\otimes_{A}\kappa_{i}$ then $\Theta_{i}\in\mathrm{cPC}^{\Gamma}_{G}(\kappa_{i})$ and $\widetilde{\Theta}_{i}=\Theta\otimes_{A}\hat{A}_{\mathfrak{m}_{i}}$ is a deformation of $\Theta_{i}$ to $\hat{A}_{\mathfrak{m}_{i}}$. 7.5 implies that $\Gamma/\operatorname{Ker}(\widetilde{\Theta}_{i})$ is an extension of a pro-$p$ group by a finite group. Hence the same applies to $\Gamma/\bigcap_{i=1}^{r}\operatorname{Ker}(\widetilde{\Theta}_{i})$. Since $A$ injects into $\prod_{i=1}^{r}\hat{A}_{\mathfrak{m}_{i}}$ we conclude that $\Gamma/\operatorname{Ker}(\Theta)$ is a quotient of $\Gamma/\bigcap_{i=1}^{r}\operatorname{Ker}(\widetilde{\Theta}_{i})$ and the assertion follows. ∎
@@
Let $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(A)$, where $A$ is an affinoid $L$-algebra. If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then there exists an $\mathcal{O}$-subalgebra $A_{0}\subseteq A$ which is a quotient of $\mathcal{O}\langle x_{1},\dots,x_{t}\rangle$ and $A_{0}[1/p]=A$, such that $\Theta$ takes values in $A_{0}$.
@@
Our assumption on $\Gamma$ and 8.1 imply that $\Gamma/\operatorname{Ker}(\Theta)$ is finitely generated. The assertion follow from [14, Lemma 6.16 (2)]. ∎
Let $|\mathrm{PC}^{\Gamma}_{G}|\subset\mathrm{PC}^{\Gamma}_{G}$ be the subset of closed points $z$ with finite residue field $k_{z}$, such that the canonical $G$-pseudocharacter $\Theta_{z}\in\mathrm{PC}^{\Gamma}_{G}(k_{z})$ attached to $z$ is continuous for the discrete topology on $k_{z}$. Let $\mathcal{O}_{z}=\mathcal{O}\otimes_{W(k)}W(k_{z})$ and let $R^{\mathrm{ps}}_{\Theta_{z}}$ be the deformation ring representing $\operatorname{Def}_{\Theta_{z}}:\mathfrak{A}_{\mathcal{O}_{z}}\rightarrow\operatorname{Set}$ and let $\Theta^{u}_{z}\in\mathrm{cPC}^{\Gamma}_{G}(R^{\mathrm{ps}}_{\Theta_{z}})$ be the universal deformation of $\Theta_{z}$. If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then $R^{\mathrm{ps}}_{\Theta_{z}}$ is a complete local noetherian $\mathcal{O}_{z}$-algebra with residue field $k_{z}$ by 6.2, and we let $(\operatorname{Spf}R^{\mathrm{ps}}_{\Theta_{z}})^{\mathrm{rig}}$ be the Berthelot rigid generic fibre of the formal scheme $\operatorname{Spf}R^{\mathrm{ps}}_{\Theta_{z}}$, [8, Section 7].
@@
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then the functor $\tilde{X}_{G}$ is representable by a quasi-Stein space $X_{G}:=\coprod_{z\in|\mathrm{PC}^{\Gamma}_{G}|}(\operatorname{Spf}R^{\mathrm{ps}}_{\Theta_{z}})^{\mathrm{rig}}$.
@@
Let $A$ be an $L$-affinoid algebra, and let $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(A)$. It follows from 8.1 that $\Theta\in\mathrm{cPC}^{Q}_{G}(A)$ for a finitely generated quotient $Q$ of $\Gamma$. Since the assertion of the theorem for finitely generated groups $\Gamma$ is proved in [14, Theorem 6.1], $\Theta$ induces a morphism of rigid analytic spaces:
@@
If $\Gamma$ is finitely generated then the assertion is proved in [14, Corollary 6.23] and the same proof carries over. By the definition of $X_{G}$, we have $X_{G}(\overline{L})=\mathrm{cPC}^{\Gamma}_{G}(\overline{L})$, where $\overline{L}$ carries the direct limit topology of its finite subextensions of $L$. The claim now follows from the reconstruction theorem [14, Theorem 3.8]. ∎

### v2

@@ 7. Completions at maximal ideals and deformation problems
Let $\kappa$ be a local field that is an $\mathcal{O}$-algebra. Let $\mathcal{O}_{\kappa}$ be its ring of integers and let $k^{\prime}$ be the residue field. We always equip $\kappa$ with the topology induced by the valuation. Let $\mathcal{O}^{\prime}=\mathcal{O}\otimes_{W(k)}W(k^{\prime})$, where $W(k)$ is the Witt ring of $k$.
In this section we will study deformation theory of $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(\kappa)$, where $\Gamma$ is an arbitrary profinite group and $G$ is a generalised reductive group scheme over $\mathcal{O}$. If $\mathrm{char}(\kappa)=0$ then we let $\Lambda=\kappa$ equipped with its natural topology and we let $\Lambda_{0}=\mathcal{O}_{\kappa}$. If $\mathrm{char}(\kappa)=p$ then [^5] The study of deformation problem in this setting is motivated by the work of Böckle–Juschka [5, Section 4.7]. we let $\Lambda=\mathcal{O}\otimes_{W(k)}C$, where $C$ is a Cohen ring of $\kappa$, which means that $C$ is a complete DVR with uniformiser $p$ and residue field $\kappa$. If we choose an isomorphism $k^{\prime}\llbracket t\rrbracket\cong\mathcal{O}_{\kappa}$ of local $k^{\prime}$-algebras then $C$ can be chosen to be the $p$-adic completion of $W(k^{\prime})\llbracket t\rrbracket[1/t]$. There is a natural topology on $\Lambda$, such that the quotient topology on $\kappa$ induced via $\Lambda\twoheadrightarrow\kappa$ coincides with the valuation topology on $\kappa$. The topology on $\Lambda$ is Hausdorff, see [4, Section 3.5] for more details. We let $\Lambda_{0}=\mathcal{O}^{\prime}\llbracket t\rrbracket$. In both cases we let $\mathfrak{A}_{\Lambda}$ be the category of local artinian $\Lambda$-algebras with residue field $\kappa$. Any $A\in\mathfrak{A}_{\Lambda}$ is a finitely generated $\Lambda$-module, and hence carries a unique Hausdorff topology, which makes $A$ into a topological $\Lambda$-module.
For $A\in\mathfrak{A}_{\Lambda}$ we let $\operatorname{Def}_{\Theta}(A)$ be the set of $\widetilde{\Theta}\in\mathrm{cPC}^{\Gamma}_{G}(A)$ such that $\widetilde{\Theta}\otimes_{A}\kappa=\Theta$. It follows [^6] If $\mathrm{char}(\kappa)=p$ then $\Lambda$ is taken to be equal to $\kappa$ in [16]. However, the argument used in the proof of [16, Theorem 5.4] goes through in our more general setting. from [16, Theorem 5.4] that the functor $\operatorname{Def}_{\Theta}:\mathfrak{A}_{\Lambda}\rightarrow\operatorname{Set}$ is pro-represented by a Hausdorff topological local $\Lambda$-algebra $R^{\mathrm{ps}}_{\Theta}$, which is a projective limit of rings in $\mathfrak{A}_{\Lambda}$. We let $\widetilde{\Theta}^{u}\in\mathrm{cPC}^{\Gamma}_{G}(R^{\mathrm{ps}}_{\Theta})$ be the the universal deformation of $\Theta$.
@@
By the reconstruction theorem [16, Theorem 3.7] there exists a continuous representation $\rho:\Gamma\rightarrow G(\overline{\kappa})$, such that $\Theta_{\rho}=\overline{\Theta}\otimes_{\kappa}\overline{\kappa}$. If $\kappa^{\prime}$ is a finite extension of $\kappa$ contained in $\overline{\kappa}$ then $G(\kappa^{\prime})$ is a closed subgroup of $G(\overline{\kappa})$ and we let $\Gamma_{\kappa^{\prime}}$ be its preimage in $\Gamma$. Then $\Gamma_{\kappa^{\prime}}$ is a closed and hence profinite subgroup of $\Gamma$. It follows from [7, Lemma 4.3] that there exists a finite extension $\kappa^{\prime\prime}$ of $\kappa^{\prime}$ contained in $\overline{\kappa}$ and $g\in G^{0}(\kappa^{\prime\prime})$ such that $g\rho(\Gamma_{\kappa^{\prime}})g^{-1}\subset G(\mathcal{O}_{\kappa^{\prime\prime}})$. This implies that $\Theta|_{\Gamma_{\kappa^{\prime}}}$ takes values in $\mathcal{O}_{\kappa^{\prime\prime}}$. Since $\Gamma=\bigcup_{\kappa^{\prime}}\Gamma_{\kappa^{\prime}}$ we deduce that $\Theta$ takes values in $\mathcal{O}_{\overline{\kappa}}\cap\kappa=\mathcal{O}_{\kappa}$. ∎
@@
Let $B$ be the $\mathfrak{q}$-adic completion of $\Lambda\otimes_{\mathcal{O}^{\prime}}R^{\mathrm{ps}}_{\overline{\Theta}}$. Then $B$ is a complete local noetherian $\Lambda$-algebra with residue field $\kappa$ by [5, Lemma 3.3.5], [4, Lemma 3.36]. Thus $B/\mathfrak{m}_{B}^{n}\in\mathfrak{A}_{\Lambda}$ for all $n\geq 1$ and $\widetilde{\Theta}_{B}:=\Theta^{u}\otimes_{R^{\mathrm{ps}}_{\overline{\Theta}}}B$ is a deformation of $\Theta$ to $B$.
The ring $B$ pro-represents a subfunctor $\operatorname{Def}^{\prime}_{\Theta}$ of $\operatorname{Def}_{\Theta}$, such that if $A\in\mathfrak{A}_{\Lambda}$, then $\widetilde{\Theta}\in\operatorname{Def}_{\Theta}(A)$ lies in $\operatorname{Def}^{\prime}_{\Theta}(A)$ if and only if $\widetilde{\Theta}$ takes values in a $\Lambda_{0}$-subalgebra $A_{0}$ of $A$, which is finite as a $\Lambda_{0}$-module. Indeed, in this case $\widetilde{\Theta}$ is a deformation of $\overline{\Theta}$ to $A_{0}$, and since $A_{0}$ is a limit of rings in $\mathfrak{A}_{\mathcal{O}^{\prime}}$ we obtain a morphism of local $\mathcal{O}^{\prime}$-algebras $R^{\mathrm{ps}}_{\overline{\Theta}}\rightarrow A_{0}$, which induces a morphism $B\rightarrow A$, such that $\widetilde{\Theta}=\widetilde{\Theta}_{B}\otimes_{B}A$. Moreover, $\widetilde{\Theta}_{B}\otimes_{B}B/\mathfrak{m}_{B}^{n}$ takes values in the image of $\Lambda_{0}\otimes_{\mathcal{O}^{\prime}}R^{\mathrm{ps}}_{\overline{\Theta}}$ in $B/\mathfrak{m}_{B}^{n}$, which is a finite $\Lambda_{0}$-module.
Since $\Gamma$ is finitely generated as a topological group, we may choose a dense discrete finitely generated subgroup $\Sigma$ of $\Gamma$. Since $\Sigma$ is dense the restriction map induces an injection $\operatorname{Def}_{\Theta}(A)\subset\mathrm{PC}^{\Sigma}_{G}(A)$. Let $A_{\infty}$ be the $\Lambda_{0}$-subalgebra of $A$ equal to the preimage of $\mathcal{O}_{\kappa}$ under $A\twoheadrightarrow\kappa$. Since $\mathfrak{m}_{A}$ nilpotent and a finite $\Lambda$-module, $A_{\infty}$ is a union of $\Lambda_{0}$-subalgebras $A_{n}$ of $A$, $n\geq 1$, such that each $A_{n}$ is a finite $\Lambda_{0}$-module and the image of $A_{n}\rightarrow\kappa$ is equal to $\mathcal{O}_{\kappa}$. 7.1 implies that the restriction to $\Sigma$ identifies $\operatorname{Def}_{\Theta}(A)$ with a subset of $\mathrm{PC}^{\Sigma}_{G}(A_{\infty})$. Since $\Sigma$ is finitely generated, $\mathrm{PC}^{\Sigma}_{G}$ is represented by a finite type $\mathcal{O}$-algebra by [16, Proposition 3.21]. Hence, $\mathrm{PC}^{\Sigma}_{G}(A_{\infty})$ is the union of $\mathrm{PC}^{\Sigma}_{G}(A_{n})$ for $n\geq 1$. This implies that $\operatorname{Def}^{\prime}_{\Theta}(A)=\operatorname{Def}_{\Theta}(A)$, which implies the assertion. ∎
As in Section 5, let $\overline{\rho}:\Gamma\rightarrow G(\bar{k})$ be a continuous, $G$-semisimple representation with $\Theta_{\overline{\rho}}=\overline{\Theta}$. Let $K$ be the kernel of $\overline{\rho}$ and let $Q$ be the quotient of $\Gamma$ defined in (12).
@@
It follows from 7.2 that $\operatorname{Ker}(\Theta^{u})\subset\operatorname{Ker}(\widetilde{\Theta}^{u})$. Since $K/\operatorname{Ker}(\Theta^{u})$ is pro-$p$ by 5.5, $K/\operatorname{Ker}(\widetilde{\Theta}^{u})$ is also pro-pp. Thus $\Gamma/\operatorname{Ker}(\widetilde{\Theta}^{u})$ is a quotient of $Q$, which implies the equality of deformation rings. ∎
@@ 8. The rigid analytic space of $G$-pseudocharacters
Let $\mathrm{Rig}_{L}$ be the category of rigid analytic spaces over $L$. If $Y\in\mathrm{Rig}_{L}$ then its ring of global sections $\mathcal{O}(Y)$ carries a natural Hausdorff topology. If $Y=\operatorname{Sp}(A)$ is affinoid then the topology on $A$ is the Banach space topology. Let $\tilde{X}_{G}:\mathrm{Rig}_{L}^{\mathrm{op}}\rightarrow\operatorname{Set}$ be the functor that associates to every rigid analytic space $Y\in\mathrm{Rig}_{L}$ the set of continuous $G$-pseudocharacters $\mathrm{cPC}^{\Gamma}_{G}(\mathcal{O}(Y))$. We show that if $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then $\tilde{X}_{G}$ is represented by a rigid analytic space.
@@
Let $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(A)$, where $A$ is an affinoid $L$-algebra. Then $\Gamma/\operatorname{Ker}(\Theta)$ is an extension of a finite group by a pro-$p$ group.
@@
Let $\kappa_{i}$ be the residue field of $\mathfrak{m}_{i}$. Since $A$ is an affinoid algebra, $\kappa_{i}$ is a finite extension of $\mathbb{Q}_{p}$. Let $\Theta_{i}=\Theta\otimes_{A}\kappa_{i}$ then $\Theta_{i}\in\mathrm{cPC}^{\Gamma}_{G}(\kappa_{i})$ and $\widetilde{\Theta}_{i}=\Theta\otimes_{A}\hat{A}_{\mathfrak{m}_{i}}$ is a deformation of $\Theta_{i}$ to $\hat{A}_{\mathfrak{m}_{i}}$. 7.5 implies that $\Gamma/\operatorname{Ker}(\widetilde{\Theta}_{i})$ is an extension of a finite group by a pro-$p$ group. Hence the same applies to $\Gamma/\bigcap_{i=1}^{r}\operatorname{Ker}(\widetilde{\Theta}_{i})$. Since $A$ injects into $\prod_{i=1}^{r}\hat{A}_{\mathfrak{m}_{i}}$ we conclude that $\Gamma/\operatorname{Ker}(\Theta)$ is a quotient of $\Gamma/\bigcap_{i=1}^{r}\operatorname{Ker}(\widetilde{\Theta}_{i})$ and the assertion follows. ∎
@@
Let $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(A)$, where $A$ is an affinoid $L$-algebra. If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then there exists an $\mathcal{O}$-subalgebra $A_{0}\subseteq A$ which is a quotient of $\mathcal{O}\langle x_{1},\dots,x_{t}\rangle$ and $A_{0}[1/p]=A$, such that $\Theta$ takes values in $A_{0}$.
@@
Our assumption on $\Gamma$ and 8.1 imply that $\Gamma/\operatorname{Ker}(\Theta)$ is finitely generated. The assertion follows from [16, Lemma 6.16 (2)]. ∎
Let $|\mathrm{PC}^{\Gamma}_{G}|\subset\mathrm{PC}^{\Gamma}_{G}$ be the subset of closed points $z$ with finite residue field $k_{z}$, such that the canonical $G$-pseudocharacter $\Theta_{z}\in\mathrm{PC}^{\Gamma}_{G}(k_{z})$ attached to $z$ is continuous for the discrete topology on $k_{z}$. Let $\mathcal{O}_{z}=\mathcal{O}\otimes_{W(k)}W(k_{z})$ and let $R^{\mathrm{ps}}_{\Theta_{z}}$ be the deformation ring representing $\operatorname{Def}_{\Theta_{z}}:\mathfrak{A}_{\mathcal{O}_{z}}\rightarrow\operatorname{Set}$ and let $\Theta^{u}_{z}\in\mathrm{cPC}^{\Gamma}_{G}(R^{\mathrm{ps}}_{\Theta_{z}})$ be the universal deformation of $\Theta_{z}$. If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$ then $R^{\mathrm{ps}}_{\Theta_{z}}$ is a complete local noetherian $\mathcal{O}_{z}$-algebra with residue field $k_{z}$ by 6.2, and we let $(\operatorname{Spf}R^{\mathrm{ps}}_{\Theta_{z}})^{\mathrm{rig}}$ be the Berthelot rigid generic fibre of the formal scheme $\operatorname{Spf}R^{\mathrm{ps}}_{\Theta_{z}}$, [9, Section 7].
@@
If $\Gamma$ satisfies Mazur’s condition $\Phi_{p}$, then the functor $\tilde{X}_{G}$ is representable by a quasi-Stein space $X_{G}:=\coprod_{z\in|\mathrm{PC}^{\Gamma}_{G}|}(\operatorname{Spf}R^{\mathrm{ps}}_{\Theta_{z}})^{\mathrm{rig}}$.
@@
Let $A$ be an $L$-affinoid algebra, and let $\Theta\in\mathrm{cPC}^{\Gamma}_{G}(A)$. It follows from 8.1 that $\Theta\in\mathrm{cPC}^{Q}_{G}(A)$ for a finitely generated quotient $Q$ of $\Gamma$. Since the assertion of the theorem for finitely generated groups $\Gamma$ is proved in [16, Theorem 6.1], $\Theta$ induces a morphism of rigid analytic spaces:
@@
If $\Gamma$ is finitely generated then the assertion is proved in [16, Corollary 6.23] and the same proof carries over. By the definition of $X_{G}$, we have $X_{G}(\overline{L})=\mathrm{cPC}^{\Gamma}_{G}(\overline{L})$, where $\overline{L}$ carries the direct limit topology of its finite subextensions of $L$. The claim now follows from the reconstruction theorem [16, Theorem 3.8]. ∎